OpenStudy (anonymous):

Whats the formula for finding the radius of a circular sector given the angle and arc length?

OpenStudy (saifoo.khan):

\[Arc.Length = \frac{x}{36}\times 2 \pi r\]

OpenStudy (anonymous):

\[L = 2\pi r \times \frac{\theta}{2\pi} = \theta r\] So rearrange for r: \[r = \frac{L}{\theta}\] Of course this only applies if the angle is in radians, degrees can be converted easily.

OpenStudy (anonymous):

arc length = radius * angle in radians radius = arc length / angle in radians

OpenStudy (anonymous):

if angle is in degrees then radians = degrees * pi /180

OpenStudy (anonymous):

okay so i do 15/(3pi/2) and i get 10pi but its wrong? what am i doing wrong?

OpenStudy (anonymous):

what is the question exactly?

OpenStudy (anonymous):

\[\frac{15}{\frac{3\pi}{2}} = \frac{30}{3\pi} = \frac{10}{\pi} \approx 3.183\]

OpenStudy (anonymous):

thats correct